Item 20
A chemist has 0.2 liters of an acid solution whose acid concentration is 65%. How many liters of pure water should be added to lower the concentration to 50%?

Answer by stanbon(75887)

Item 20
A chemist has 0.2 liters of an acid solution whose acid concentration is 65%. How many liters of pure water should be added to lower the concentration to 50%?
----------------
Equation:
acid + acid = acid
0.65*0.2 + 0*x = 0.50(0.2+x)
------------------------------
65*0.2 + 0 = 50*0.2 + 50x
15*0.2 = 50x
x = 0.3*0.2
x = 0.06 liters (amt of pure water to add)
-------------
Cheers,
Stan H.
----------